OpenStudy (anonymous):

If log x = 1/2 log a - log b and a = 4b^2, then x = (a) 1 (b) 2 (c) 4 (d) 8 (e) 16

OpenStudy (anonymous):

ok let a=4b^2 in the above log x=1/2log (4b^2)-log b or using property of log \[\Large \log x=\log (4b^2)^{\frac{1}{2}}-\log b\] \[\Large \log x=\log(2b)-\log(b)\] now again using log property \[\Large \log x=\log(2b/b)\] \[\Large \log x= \log 2\]

OpenStudy (anonymous):

@waterineyes do not you think it should be x=2 ??

OpenStudy (anonymous):

\[\log(x) = \frac{1}{2}\log(2b)^2 - \log(b) \implies \log(x) = \frac{1}{2} \times 2 \log(2b) - \log(b)\] \[\log(x) = \log(2b) - \log(b) = \implies \log(x) = \log(2) \implies \color{blue}{ x = 2}\]

OpenStudy (anonymous):

Yeah sorry my mistake there @sami-21

OpenStudy (anonymous):

its ok :) i do a lot of mistakes . i am notorious for typo mistakes here :P ask anyone here :P

OpenStudy (anonymous):

@lawls understood or any problem anywhere ??

OpenStudy (anonymous):

hold on i'm reading what u guys wrote

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

ok i'm done reading. thanks alot guys i fully understand now :)

OpenStudy (anonymous):

Sure ??

OpenStudy (anonymous):

yeah x is 2 right?

OpenStudy (anonymous):

yes x=2